Part A A student measures the OH- concentration in a saturated aqueous solution of nickel(II) hydroxide...

Part A

A student measures the OH- concentration in a saturated aqueous solution of nickel(II) hydroxide to be 8.44×10-6 M. Based on her data, the solubility product constant for nickel(II) hydroxide is

Part B

A student measures the Pb2+ concentration in a saturated aqueous solution of lead bromide to be 1.14×10-2 M.
Based on her data, the solubility product constant for lead bromide is

Part C

A student measures the molar solubility of silver carbonate in a water solution to be 1.29×10-4 M.
Based on her data, the solubility product constant for this compound is
Answer #1

Part A)

Ni(OH)2 solubilizes as

Ni(OH)2\rightleftharpoons Ni2+(aq) + 2 OH-(aq)

                    x              2x

Since OH- is 8.44×10-6 M Ni2+ will be 8.44×10-6 M/2 = 4.22 ×10-6 M

Ksp = [Ni2+][OH-]2

Ksp = 4.22 ×10-6 x (8.44×10-6)2

Ksp = 3.00 x 10-16

Part B

A student measures the Pb2+ concentration in a saturated aqueous solution of lead bromide to be 1.14×10-2 M.

PbBr2\rightleftharpoons Pb2+(aq) + 2 Br-(aq)

                    x              2x

Since Pb2+ is 1.14×10-2 M Br- is 1.14×10-2 M x 2 = 2.28 ×10-2 M

Ksp = [Pb2+][Br-]2

Ksp = 1.14×10-2 x (2.28 ×10-2)2

Ksp = 5.92 x 10-6

Part C

A student measures the molar solubility of silver carbonate in a water solution to be 1.29×10-4 M.

Ag2CO3\rightleftharpoons 2Ag+ (aq) + CO32-(aq)

molar solubility is 1.29×10-4 M so silver ions are 2 x 1.29×10-4 M = 2.58 ×10-4 M

Ksp = [Ag+]2[CO32-]

= (2.58 ×10-4)2 (1.29×10-4)

Ksp = 8.58 x 10-12